Status Report

ISS On-Orbit Status 26 Feb 2002

By SpaceRef Editor
February 26, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

All activity on R2 (release 2) software upgrading for U.S. segment control computers have been successfully completed. Several procedural problems had to be overcome, but overall the transition went very well. CSA (Canadian Space Agency) reports that the SSRMS/Canadarm2 is working nominally after its R2 s/w upgrade, which provides arm ops with new capabilities.

Crew today completed a familiarization session with the robotics DOUG (Dynamic Onboard Ubiquitous Graphics) application on their standalone PCS (portable computer system), preparatory to tomorrow’s “Robotics Day #2”. DOUG is a s/w program that provides a birdseye-view graphical image of the SSRMS arm, showing its real-time location and configuration on a laptop during its operation. [Robotics Day #2 will be dedicated to complete some exterior surveys of ISS, specifically the radiator array panels and base plates for any surface discolorations or deformations (bubbles, flakes) as small as 1 cm. The surveys will be videotaped for later downlink and the tape returned on 8A.]

The crew went through one of the three-hour Emergency Descent OBT (onboard training) drills which are scheduled for each Expedition, to refresh crew safety/survival acuity in case of an event requiring early station evacuation. For the sim, the crew reviewed applicable FDF (flight data file) sections, going through all necessary steps from donning the Sokol entry suits inside the Soyuz to automatic undock and descent, even considering a leaky Sokol. They also practiced Soyuz descent module Manual Entry Control (RUS) mode using an on-board trainer program. The necessary software to take them through the OBT is set up prior to the sim on an SSC (space station computer) laptop.

CDR Onufrienko completed regular daily servicing tasks, i.e., switching regeneration cycle from filter #1 of the BMP micropurification unit to filter #2, maintenance of SOSH (Russian ECLSS) life support system and inspection of the BRPK-1 condensate separation and pumping unit [there are two of these systems in the condensate water processor, but only one is operating at a time, removing condensate from the gas-liquid mixture and pumping it to the purification columns unit (BKO) and water conditioning unit (BKV)].

FE-1 Walz prepared today’s IMS (inventory management system) update (“delta”) file and checked on Increment 4 payload status in the Lab.

FE-2 Bursch took historic photography of the ADVASC (advanced astroculture) experiment, focusing on a close-up of the experiment set-up from the front, with growth chamber and support system in the same shot, and on the entire EXPRESS Rack with ADSVASC installed in it.

SD2 (smoke detector #2) in the Node was reported failed two days ago. While troubleshooting is underway on the ground, the unit is kept inhibited. SD1 continues to function and remains enabled.

At about 9:00am EST, MCC-H started reinitialization of the two external batteries of the P6 4B solar array wing. The activity will take about two days to complete. Ê[Reason: There is a significant (and increasing) internal pressure difference between the two batteries in the set, which results in uneven charging on the two storage units. To reinitialize, the two batteries in will be completely discharged down to empty, and then fully recharged, which should get them back at the same pressure level. With the current load carried on channel 4B, no power-downs are required during this activity. There are no crew actions for this, but several suppressed caution alerts are expected.]

Late last evening, CMG-3 (control moment gyroscope #3) momentarily lost communications with the electronic assembly associated with it. Later during the night, the same happened with CMG-4. This temporary drop-out has been seen on all four CMGs since they were brought on line, but the cause has not yet been isolated (suspicion is on a timing problem). There is automatic software in place which brings the CMG back into the control loop within two minutes, and this also worked this time. The U.S. segment’s MCS (motion control system) is running nominally.

His morning the TVIS (treadmill with vibration isolation and stabilization) experienced an apparent power failure, slowing down and stopping while Carl Walz was using it. Some limited troubleshooting did not regain its capability, and a more thorough plan is being developed on the ground. If TVIS should remain down for longer, MedOps personnel have alternative exercise programs for providing the crew with the all-important aerobics workout.

Regeneration of the partially regenerated Metox (metal oxide) canisters has been postponed four days. [Of the four Metox canisters used during 7A last July, the two EMU suit canisters had been properly regenerated. The bake-out of the two remaining cans, used to scrub CO2 and other contaminants from the Joint Airlock (JAL) atmosphere, had been terminated prematurely to investigate a noise caused by the air conditioning system. After being left in the oven for seven months, they were heated again on 2/20 after the EVA by Bursch and Walz, generating disturbing odors]. Before their regeneration can be completed, concurrence is required by the joint (US/Russian) ECLSS and Safety teams. Their recommendations will be reviewed at next Friday’s MMT (Mission Management Team) meeting, and the procedure, if approved, will be scheduled on Saturday, 3/2 (Monday, 3/4, as backup day). It is not expected at this time that new Metox canisters have to be brought up by 8A.

A troubleshooting plan for a bad hard drive bay for the EXPPCS experiment was uplinked to restore the science results for this payload as soon as possible. The plan essentially removes two full hard drives and relocates the operating system hard drive to one of the freed bays.

The crew, looking great, participated in two live media interviews, starting at 11:05 am EST. One was with WICZ-TV of Vestal, NY, the other with WOIO-TV of Cleveland, OH, each for about 10 minutes.

The solar Beta angle has reached its maximum overnight and is now starting to decrease again.

An EVA debriefing with the crew is being scheduled for Thursday morning, 2/28 (day of launch of STS-109).

Launch date of Progress 7P (M1-257) continues to be 3/21 (backup day: 3/22). MCC-M/TsUP confirmed that two orbit correction/reboost maneuvers are scheduled for 3/6 and 3/13, using the eight small thrusters of Progress 6P (M1-256), with propellants pumped over from the FGB.

Today’s targets for CEO (crew earth observations) were E. Mediterranean Dust and Smog (with high pressure building into the eastern Mediterranean this pass, crew was to look to the left of track as it crossed the Libyan coast toward Crete and Turkey for the development of smog layers in the atmosphere), W. Mediterranean Dust and Smog (fair weather and low light this pass favored views of aerosol plumes over the darker sea surfaces. Of interest: looking to the left of track over both the west and east coasts of Italy in oblique views), Eastern United States (cold high pressure persists over the Eastern US. This pass, looking to the right of track over the Appalachians may have shown evidence of aerosol accumulations in oblique views. The second pass should have been an excellent pass for detecting aerosols with oblique and limb views over the upper Ohio River valley and the northeastern US. As crew approached Lake Erie from the NW, it was to look to the left of track), European Smog (pass paralleled the western coast of Italy. After crossing the French Rivera, crew was to look to the left of track and document smog concentrations and patterns over Italy and its adjacent coastal waters).

U.S. and Russian Segment Status (as of 1:45 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16 Amp mode, the lowest possible setting). Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.

  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Regeneration.

  • SM Working Compartment: Pressure (mmHg) — 764, temperature (deg C) — 26.5, ppO2 (mmHg) — 160.0, ppCO2 (mmHg) — 2.6 (? see Notes below and above)

  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) 21.5; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — n/a; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • Node: Pressure (mmHg) — 757.25, temperature (deg C) — n/a (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • U.S. Lab: Pressure (mmHg) — 759.06, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;

  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 758.75, temperature (deg C) — n/a; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

  • PMA-1: Shell heater temp (deg C) — n/a.

  • PMA-2: Shell heater temp (deg C) — n/a.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]. Russian GA (gas analyzer) readings of ppCO2 in the SM are invalid).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (during XPOP).

  • SM batteries: all batteries (8) in “partial charge” mode.

  • FGB battery #5 is offline; all other batteries (5) are in “partial charge” mode.

  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-1 is back-up, and C&C-2 is in standby, all with new R2 s/w.

  • GNC-2 MDM is prime; GNC-1 is back-up, with R1 s/w.

  • LA-1, LA-2 and LA-3 MDMs are all operating.

  • APS-1 (automated payload switch #1) and APS-2 are both On.

  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.

  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.

  • Ku-band is operating nominally in open loop pointing mode.

  • Audio subsystem operating nominally.

  • Video subsystem operating nominally.

  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.

  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 7:43am EST):

  • Mean altitude — 384.2 km

  • Apogee — 391.5 km

  • Perigee — 376.8 km

  • Period — 92.2 min.

  • Inclination (to Equator) — 51.64 deg

  • Eccentricity — 0.0010901

  • Orbits per 24-hr. day — 15.61

  • Altitude decrease — 315 m (mean) in last 24 hours

  • Solar Beta Angle: -60.6 deg (magnitude decreasing)

  • Revolutions since FGB/Zarya launch (Nov. ’98) — 18677

  • Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane [yaw: ~0 deg, pitch: 5 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.